• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name B. Craver

    Log entry time 22:45:24 on March 5,2007

    Entry number 194897

    keyword=beam spot

    For run 1193, with the beam moved -1mm in the x position relative to the production runs 1188 and 1189, the position of the beam looks more uniform. The raster currents are still non-uniform, but the current hypothesis is that the beam was scraping on the edge and now is not. T3 has dropped a small amount to ~5 kHz for the current run from ~6 kHz previously.


    FIGURE 1

    FIGURE 2